OpenStudy (anonymous):

1/(y-1)(y-2) = A/(y-1)+B/(y-2) How to do you find A and B?

OpenStudy (anonymous):

there is two why simple one and exact one

OpenStudy (anonymous):

i need the exact one, i suppose

OpenStudy (anonymous):

let me explain them for you : the exact one tells multiplay second side and compare nominator of both side.

OpenStudy (anonymous):

i mean : a(y-2)+b(y-1)= (a+b)y-(2a+b)

OpenStudy (anonymous):

know you can cancel both side dominator with each other .

OpenStudy (anonymous):

now compare what remains, coefficient of y with same degree should be equal .

OpenStudy (anonymous):

look at your first equation you just have 1 . it 's equal with 0*y + 1

OpenStudy (anonymous):

so 0 = a+b -2a-b=1 a=-1 b=1

OpenStudy (anonymous):

if you get this way i could tell you the simple way

OpenStudy (anonymous):

ok i'm trying to understand...

OpenStudy (anonymous):

if u have any problem you could ask

OpenStudy (anonymous):

a(y-2)+b(y-1)= (a+b)y-(2a+b) how do you get the right part (after the = sign)?

OpenStudy (anonymous):

oh no it's ok, i got it...

OpenStudy (anonymous):

sorry i was busy with another question

OpenStudy (anonymous):

what do you mean by the coefficient of y should be equal?

OpenStudy (phi):

the idea is compare (for example) 2y^2 + 3y + 1 = ay^2 + by + c in this example, for the two sides to be equal (for all y), it requires that a is 2, b is 3 and c is 1 use that same idea in your problem.

OpenStudy (phi):

in your problem you have 1 = (a+b)y-(2a+b)

OpenStudy (anonymous):

i don't know the exact word in english but phi get what i mean

OpenStudy (phi):

or 0y + 1 = (a+b)y-(2a+b)

OpenStudy (anonymous):

why do you put 0y? for what purpose?

OpenStudy (phi):

we started with 1/(y-1)(y-2) = A/(y-1)+B/(y-2) and (after getting a common denominator) \[ \frac{1}{(y-1)(y-2)} = \frac{(A+B) y - (2A+B)}{(y-1)(y-2)}\] equate the numerators \[ 1=(A+B) y - (2A+B) \] or because you have a y term on the right side, explicitly show a y term on the left side \[ 0y+ 1=(A+B) y - (2A+B) \]

OpenStudy (phi):

now you can "match coefficients"

OpenStudy (anonymous):

@ phi : thanks you tough me a new word

OpenStudy (anonymous):

so A+B = 0 and -(2A+B) = 1

OpenStudy (phi):

yes, now you have to solve 2 simultaneous equations

OpenStudy (anonymous):

and now its easy thanx a lot

OpenStudy (anonymous):

A=1 and B=-1

OpenStudy (anonymous):

A=-1 and B =1

OpenStudy (anonymous):

oh yes right

OpenStudy (anonymous):

actually no i get B=-1

OpenStudy (anonymous):

that's ur answer so A+B = 0 and -(2A+B) = 1 a=-b 2a+b=-1 2a-a=-1 a=-1 a=-b b=1

OpenStudy (anonymous):

oh yeah sry